# Score-based diffusion nowcasting of GOES imagery

### v1

@@ 5 Summary and Conclusions
The diffusion models we showed here are far from perfect. The Diff model tends to be noisier than the CorrDiff approach, which could be detrimental to longer forecast lead times. Furthermore, all machine learning methods had a clear bias with time, some trending warm, others trending cold, which could eventually lead to nonphysical forecasts. Another challenge of the diffusion models is their increased computational cost compared to the more traditional U-Net (i.e., MSE loss). It took about five days to train the plain diffusion model and the CorrDiff model on a relatively new GPU system (e.g., GH-200). This length of training time is prohibitively long to allow for a proper hyperparameter search of the underlying architecture on our resources. Furthermore, unlike in more traditional machine learning methods, the computational cost is not isolated to training time. Because of the iterative steps of denoising, the diffusion models take substantially longer to run during inference time, 10 mins to generate a three hour forecast with 10-members for the diffusion model versus the 10 seconds for the U-Net. This added computational time also prohibited hyperparameter tuning of generation parameters. Thus, the choices we made are likely sub-optimal. We also explored the use of pretrained autoencoders to enable the use of latent diffusion, which substantially improved computation costs (one day to train, one min to run forecast), but results are clearly worse than for the other diffusion models. The LDM approach might improve if a specialized autoencoder is trained for satellite imagery.
@@
This paper closely followed the methodology from Karras et al. (2022) which was also used in Mardani et al. (2025), Manshausen et al. (2024) and Price et al. (2025). Since the original Karras et al. (2022) work, there have been improvements. The improvements are mainly centered on carefully choosing the machine learning architecture (Karras et al., 2024b) and having better image generation (Karras et al., 2024a). Thus, there are improvements to the methodology directly that could improve results here. Furthermore there is work being done to lower the computational cost of diffusion models (e.g., Xu et al., 2025).
@@
Outside of changes to the methodology, other ways to improve the satellite nowcasting would be to include additional information (i.e., more conditions) to the model input. Specifically, including simple input fields like the solar zenith angle and topography should help inform the model where in the image should be warming and what the surface type is (e.g., land vs. ocean), respectively. One could simplify the task further by breaking the task into two parts, advection and alteration. Advection could be handled well by known tools like optical flow and the diffusion model could handle the generation, growth, and/or decay of clouds. In addition to the solar zenith angle and the topography, this work could be extended to include more channels of satellite imagery. This idea of multitask learning might also help because if a multi-channel diffusion model was trained using all thermal bands, then the model could potentially learn more of the 3-dimensional structure of the atmosphere, as well as the water vapor content, which are both critical for the evolution of clouds. However, including additional input and output channels will likely take careful implementation given the already computationally intensive diffusion task.
@@
Overall, we are encouraged by the prospect of score-based diffusion models for machine learning tasks in the Earth and Environmental Sciences. They provide often high-resolution features that appear more realistic than for the more traditional U-Net approaches of the past. Furthermore, the out-of-the-box uncertainty quantification offered from ensemble generation is an added benefit, often improving pixel-wise metrics while still having the information content of the individual members.

### v2

@@ 5.2 Limitations and Future Directions
The diffusion models we showed here are far from perfect. The Diff model tends to be noisier than the CorrDiff approach, which could be detrimental to longer forecast lead times. Furthermore, all machine learning methods had a clear bias with time, some trending warm, others trending cold, which could eventually lead to nonphysical forecasts. Another challenge of the diffusion models is their increased computational cost compared to the more traditional U-Net (i.e., MSE loss). It took about five days to train the plain diffusion model and the CorrDiff model on a relatively new GPU system (e.g., GH-200). This length of training time is prohibitively long to allow for a hyperparameter search of the underlying architecture on our resources. Furthermore, unlike in more traditional machine learning methods, the computational cost is not isolated to training time. Because of the iterative steps of denoising, the diffusion models take substantially longer to run during inference time, 10 mins to generate a three hour forecast with 10-members for the diffusion model versus the 10 seconds for the U-Net. This added computational time also prohibited hyperparameter tuning of generation parameters. Thus, the choices we made are likely sub-optimal. We also explored the use of pretrained autoencoders to enable the use of latent diffusion, which substantially improved computation costs (one day to train, one min to run forecast), but results are clearly worse than for the other diffusion models. The LDM approach might improve if a specialized autoencoder is trained for satellite imagery.
Since we were compute constrained, we only had computational time to investigate a single common neural network architecture across all experiments. This is a key limitation of this paper. Other architectures should have been explored, including simpler ones (e.g., a U-Net with less parameters or no attention layers) that could have sped up the training and inference speed as well as more advanced networks (e.g., vision transformers and graphs). Another limitation of this work was the lack of comparison of the trained machine learning methods here to other methods of forecasting satellite brightness temperatures. Future work would ideally include forecasts from both traditional numerical weather prediction (e.g., Griffin2017; Griffin2024) and optical flow (e.g., Shakya2019). Even though the diffusion models trained here have considerable compute costs, they are likely faster than traditional NWP thus providing added benefit. A clearly outlined study comparing the three would be greatly beneficial to the community and is a topic of future research.
@@
This paper closely followed the methodology from Karras2022 which was also used in Mardani2025, Manshausen2024 and Price2025. Since the original Karras2022 work, there have been improvements. The improvements are mainly centered on carefully choosing the machine learning architecture (Karras2024a) and having better image generation (Karras2024b). Thus, there are improvements to the methodology directly that could improve results here.
@@
Outside of changes to the methodology, other ways to improve the satellite nowcasting would be to include additional information (i.e., more conditions) to the model input. Specifically, including simple input fields like the solar zenith angle and topography should help inform the model where in the image should be warming and what the surface type is (e.g., land vs. ocean), respectively. One could simplify the task further by breaking the task into two parts, advection and alteration. Advection could be handled well by known tools like optical flow and the diffusion model could handle the generation, growth, and/or decay of clouds. In addition to the solar zenith angle and the topography, this work could be extended to include more channels of satellite imagery. This idea of multitask learning might also help because if a multi-channel diffusion model was trained using all thermal bands, then the model could potentially learn more of the 3-dimensional structure of the atmosphere, as well as the water vapor content, which are both critical for the evolution of clouds.
6 Conclusions
A challenge with many machine learning methods is the prevalence of blurry solutions as a result of mean squared error losses that then have limited usefulness in human-in-the-loop applications (i.e., weather forecasters). In the literature, diffusion models suggest to be a solution to these blurry outputs while also providing useful uncertainty quantification through ensembles. For this paper, we delve into the intuition behind the success of diffusion models for providing sharper images, we explore the use of score-based diffusion models for a basic forecasting task (in our case to nowcast geostationary satellite imagery), and we highlight lessons learned from our efforts for both research applications and operational tasks.
This work has three main contributions aimed towards aiding the community in adopting diffusion modeling. First, we provide an intuitive introduction to the methodology of score-based diffusion models, discuss how to train and use score-based diffusion models, and detail often overlooked steps needed for earth science applications that are not addressed in the computer science literature (Section 2). Second, we explore various types of score-based diffusion models and evaluate how they perform for our nowcasting task of forecasting GOES channel 13 infrared brightness temperatures out to three hours. More specifically, we explore three different diffusion modeling approaches: a standard score-based diffusion model (Diff) that follows Karras2022, a residual corrective diffusion model (CorrDiff) that follows Mardani2025, and a latent diffusion model (LDM) inspired by Leinonen2023. We showed that the diffusion models were able to nowcast satellite imagery qualitatively and quantitatively better than both persistence forecasts and a mean-squared-error-trained U-Net (Figures 10-14). The best performing diffusion model was the CorrDiff approach followed by the Diff model and then the LDM. All diffusion models were able to not only advect the current clouds in the scene but also generate and decay clouds, a key weakness of optical flow techniques. We also showed that all methods were able to generate convection (i.e., localized regions of very cold brightness temperatures) despite only having two images (20 mins) of 10 micron infrared imagery and no information on the thermodynamic state of the atmosphere, the solar input, or the land surface type. Third, we provide valuable perspectives on utilizing diffusion models both for research and operationally, discussing trade-offs that need to be considered between accuracy and computing resources (Section 5).
@@ 6 Conclusions
Overall, we are encouraged by the prospect of score-based diffusion models for machine learning tasks in the Earth System sciences. They provide often high-resolution features that appear more realistic than the more traditional U-Net approaches. Furthermore, the out-of-the-box uncertainty quantification offered from ensemble generation is an added benefit, often improving pixel-wise metrics while still having the information content of the individual members. We hope that this work provides intuition on how diffusion modeling works as well as valuable tips to accelerating their use in the Earth System sciences.